OpenStudy (anonymous):

Identify the x and y intercepts of the equation 2x-3y=9

OpenStudy (anonymous):

(4.5, 0) (0, 6)

OpenStudy (anonymous):

When I worked it out I got x=4.5 y=-3

OpenStudy (anonymous):

2x-3y=9 x-intercept: x=0 -3y=9 -y=-3 y-intercept: y=0 2x=9 x=4.5

OpenStudy (anonymous):

@franciscanmonk Thanks

OpenStudy (anonymous):

you have to set it equal to 0

OpenStudy (anonymous):

yo i need that metal man please

OpenStudy (anonymous):

no probs @19 thejoker1177